# CursorCore: Assist Programming through Aligning Anything

### v2

@@ Appendix I Additional evaluation results on APEval
Appendix J Chat template
Our model’s chat template OpenAI (2023) is adapted from the ChatML template, where each message in the conversation is restricted to one of the following roles: system, history, current, user, or assistant. The assistant’s output includes both code modifications and chat interaction with the user. To indicate code changes, we use two special tokens “<|next_start|>” and “<|next_end|>” to wrap the code modification parts. This approach models Assistant-Conversation effectively and is compatible with standard ChatML templates and chatbot applications. Figure 17 illustrates an example of our chat template, while Figure 18 presents examples of the chat template when using the LC and SR modes described in Appendix B.
@@
Appendix K Prompts for evaluation
@@
Appendix L Prompts for data collection
We design specific system prompts and few-shot examples to collect high-quality training data, as we find that many examples are very difficult to complete with current LLMs, and only a few of them can be successfully completed using rough prompts. For AIprogrammer, we utilize LLMs to simulate programmers at three different skill levels, with each level using a distinct set of prompts as shown in Tables 7, 8 and 9. Additionally, prompts used for evaluating whether the outputs align with user intent, generating user instructions, and facilitating chat interactions between models and users are outlined in Tables 12, 10 and 11. Partial few-shot examples are shown in Figures 19, 20, 21, 22, 23 and 24.
Appendix M Limitations and future work
@@
In this work, we focus on supporting the development of single files or function-level code. However, real-world development operates at the repository level, involving multiple files and greater interaction with IDEs. Previous research has made notable advances in repository-level tasks such as code completion Zhang et al. (2023), issue fixing Jimenez et al. (2024), and documentation generation Luo et al. (2024a). Repository-level code assistance deals with larger datasets, and achieving optimal performance and speed will require more effort. We leave the exploration of multi-file repository-level programming assistance and leveraging additional IDE interactions for future work.
@@
We have only tested our models’ code assistance capabilities on Python-specific benchmarks. While multi-language program synthesis benchmarks like Multipl-E Cassano et al. (2023a) can evaluate coding abilities across languages, dedicated benchmarks are still needed to assess programming assistance for each language. Additionally, our benchmark is relatively small and based on an extension of HumanEval, making it insufficient to cover all development scenarios. Beyond using the classic Pass@k metric to evaluate accuracy, other criteria should also be considered, such as evaluating the model’s efficiency, security, and redundancy Huang et al. (2024); Pearce et al. (2021); Li et al. (2024).
@@
Methods like PPO Schulman et al. (2017) and DPO Rafailov et al. (2023), which optimize models based on human preferences, have been widely used in LLMs. In programming assistance, programmers can provide feedback on predicted outputs for identical or similar coding processes, further optimizing the model Shinn et al. (2023). To enable this, a significant amount of feedback data from programmers using AI-assisted tools should be collected or synthesized.
@@
We aim to integrate function calls Patil et al. (2023) into the model to further enhance its capabilities. One potential application is incorporating function calls into the thinking process, such as retrieving information or executing partial code for feedback. Although our final models excludes this thinking step due to performance and speed considerations, we are exploring hybrid approaches to introduce this process while maintaining speed and combine it with other strategies for searching how to edit. Another application is leveraging function calls in output, where calling a Python script for tasks like variable replacement might be more efficient than manually generating code blocks or search-and-replace strategies. For repository-level changes, using terminal commands or IDE APIs could sometimes be a more convenient solution.

### v3

@@ Appendix I Conversation retrieval for Assistant-Conversation
Qwen2.5-72B 63.4 (61.0) 36.6 (34.1) 75.6 (63.4) 39.0 (34.1) 53.7 (48.2)
GPT-4o-Mini 17.1 (17.1) 36.6 (31.7) 78.0 (70.7) 53.7 (43.9) 46.3 (40.9)
Table 8: Evaluation results of LLMs on the C++ version of APEval.
Model C H, C C, U H, C, U Avg.
6B+ Models
DS-Coder-6.7B-Base 41.5 36.6 58.5 31.7 42.1
DS-Coder-6.7B-Inst 29.3 34.1 61.0 39.0 40.9
Yi-Coder-9B 29.3 43.9 34.1 29.3 34.1
Yi-Coder-9B-Chat 34.1 36.6 56.1 39.0 41.5
Qwen2.5-Coder-7B 39.0 41.5 65.9 39.0 46.3
Qwen2.5-Coder-7B-Inst 12.2 31.7 63.4 41.5 37.2
CursorCore-DS-6.7B 48.8 43.9 65.9 34.1 48.2
CursorCore-Yi-9B 53.7 46.3 68.3 39.0 51.8
CursorCore-QW2.5-7B 46.3 43.9 63.4 43.9 49.4
1B+ Models
DS-Coder-1.3B-Base 29.3 4.9 34.1 9.8 19.5
DS-Coder-1.3B-Inst 31.7 39.0 41.5 36.6 37.2
Yi-Coder-1.5B-Chat 24.4 26.8 17.1 22.0 22.6
Yi-Coder-1.5B 19.5 22.0 14.6 14.6 17.7
Qwen2.5-Coder-1.5B 34.1 31.7 48.8 36.6 37.8
Qwen2.5-Coder-1.5B-Inst 17.1 31.7 34.1 34.1 29.3
CursorCore-DS-1.3B 39.0 39.0 56.1 39.0 42.7
CursorCore-Yi-1.5B 46.3 34.1 56.1 34.1 42.7
CursorCore-QW2.5-1.5B 41.5 39.0 58.5 36.6 43.9
Table 9: Evaluation results of LLMs on the Java version ofAPEval.
Model C H, C C, U H, C, U Avg.
6B+ Models
DS-Coder-6.7B-Base 51.2 43.9 61.0 46.3 50.6
DS-Coder-6.7B-Inst 48.8 34.1 68.3 51.2 50.6
Yi-Coder-9B 43.9 46.3 36.6 41.5 42.1
Yi-Coder-9B-Chat 43.9 41.5 56.1 39.0 45.1
Qwen2.5-Coder-7B 63.4 53.7 70.7 46.3 58.5
Qwen2.5-Coder-7B-Inst 22.0 46.3 70.7 51.2 47.6
CursorCore-DS-6.7B 63.4 56.1 65.9 48.8 58.5
CursorCore-Yi-9B 56.1 61.0 68.3 46.3 57.9
CursorCore-QW2.5-7B 65.9 58.5 78.0 51.2 63.4
1B+ Models
DS-Coder-1.3B-Base 29.3 24.4 36.6 26.8 29.3
DS-Coder-1.3B-Inst 39.0 39.0 48.8 43.9 42.7
Yi-Coder-1.5B 24.4 29.3 24.4 36.6 28.7
Yi-Coder-1.5B-Chat 31.7 24.4 12.2 26.8 23.8
Qwen2.5-Coder-1.5B 39.0 43.9 53.7 48.8 46.3
Qwen2.5-Coder-1.5B-Inst 34.1 36.6 53.7 48.8 43.3
CursorCore-DS-1.3B 43.9 43.9 58.5 43.9 47.6
CursorCore-Yi-1.5B 53.7 48.8 53.7 46.3 50.6
CursorCore-QW2.5-1.5B 53.7 51.2 53.7 51.2 52.4
Table 10: Evaluation results of LLMs on the JavaScript version of APEval.
Model C H, C C, U H, C, U Avg.
6B+ Models
DS-Coder-6.7B-Base 29.3 26.8 51.2 39.0 36.6
DS-Coder-6.7B-Inst 48.8 41.5 58.5 36.6 46.3
Yi-Coder-9B 22.0 41.5 17.1 39.0 29.9
Yi-Coder-9B-Chat 48.8 34.1 56.1 41.5 45.1
Qwen2.5-Coder-7B 51.2 36.6 68.3 43.9 50.0
Qwen2.5-Coder-7B-Inst 22.0 34.8 63.4 43.9 40.9
CursorCore-DS-6.7B 48.8 41.5 61.0 36.6 47.0
CursorCore-Yi-9B 56.1 51.2 70.7 43.9 55.5
CursorCore-QW2.5-7B 51.2 36.6 70.7 46.3 51.2
1B+ Models
DS-Coder-1.3B-Base 19.5 9.8 26.8 26.8 20.7
DS-Coder-1.3B-Inst 29.3 34.1 36.6 36.6 34.1
Yi-Coder-1.5B 19.5 17.1 7.3 22.0 16.5
Yi-Coder-1.5B-Chat 17.1 19.5 9.8 12.2 14.6
Qwen2.5-Coder-1.5B 31.7 31.7 48.8 39.0 37.8
Qwen2.5-Coder-1.5B-Inst 19.5 26.8 43.9 39.0 32.3
CursorCore-DS-1.3B 34.1 36.6 41.5 43.9 39.0
CursorCore-Yi-1.5B 34.1 34.1 41.5 41.5 37.8
CursorCore-QW2.5-1.5B 34.1 34.1 51.2 41.5 40.2
Table 11: Evaluation results of LLMs on the Go version of APEval.
Model C H, C C, U H, C, U Avg.
6B+ Models
DS-Coder-6.7B-Base 26.8 24.4 31.7 39.0 30.5
DS-Coder-6.7B-Inst 29.3 39.0 56.1 43.9 42.1
Yi-Coder-9B 39.0 29.3 39.0 36.6 36.0
Yi-Coder-9B-Chat 34.1 24.4 41.5 31.7 32.9
Qwen2.5-Coder-7B 41.5 36.6 53.7 41.5 43.3
Qwen2.5-Coder-7B-Inst 12.2 22.0 58.5 46.3 34.8
CursorCore-DS-6.7B 56.1 34.1 56.1 46.3 48.2
CursorCore-Yi-9B 48.8 34.1 56.1 36.6 43.9
CursorCore-QW2.5-7B 58.5 41.5 58.5 51.2 52.4
1B+ Models
DS-Coder-1.3B-Base 22.0 7.3 26.8 29.3 21.3
DS-Coder-1.3B-Inst 31.7 22.0 36.6 34.1 31.1
Yi-Coder-1.5B 7.3 14.6 2.4 9.8 8.5
Yi-Coder-1.5B-Chat 19.5 14.6 17.1 29.3 20.1
Qwen2.5-Coder-1.5B 34.1 24.4 39.0 36.6 33.5
Qwen2.5-Coder-1.5B-Inst 22.0 22.0 36.6 34.1 28.7
CursorCore-DS-1.3B 41.5 34.1 43.9 41.5 40.2
CursorCore-Yi-1.5B 46.3 29.3 43.9 34.1 38.4
CursorCore-QW2.5-1.5B 51.2 34.1 61.0 41.5 47.0
Table 12: Evaluation results of LLMs on the Rust version of APEval.
Model C H, C C, U H, C, U Avg.
6B+ Models
DS-Coder-6.7B-Base 26.8 26.8 31.7 39.0 31.1
DS-Coder-6.7B-Inst 29.3 31.7 34.1 39.0 33.5
Yi-Coder-9B 29.3 34.1 29.3 39.0 32.9
Yi-Coder-9B-Chat 26.8 29.3 46.3 31.7 33.5
Qwen2.5-Coder-7B 46.3 34.1 48.8 36.6 41.5
Qwen2.5-Coder-7B-Inst 9.8 22.0 51.2 36.6 29.9
CursorCore-DS-6.7B 36.6 34.1 34.1 39.0 36.0
CursorCore-Yi-9B 41.5 34.1 48.8 41.5 41.5
CursorCore-QW2.5-7B 48.8 39.0 51.2 39.0 44.5
1B+ Models
DS-Coder-1.3B-Base 22.0 22.0 31.7 29.3 26.2
DS-Coder-1.3B-Inst 19.5 26.8 36.6 31.7 28.7
Yi-Coder-1.5B 22.0 24.4 17.1 29.3 23.2
Yi-Coder-1.5B-Chat 14.6 22.0 12.2 29.3 19.5
Qwen2.5-Coder-1.5B 34.1 29.3 39.0 41.5 36.0
Qwen2.5-Coder-1.5B-Inst 26.8 24.4 36.6 36.6 31.1
CursorCore-DS-1.3B 24.4 29.3 36.6 34.1 31.1
CursorCore-Yi-1.5B 24.4 31.7 41.5 31.7 31.1
CursorCore-QW2.5-1.5B 36.6 31.7 39.0 43.9 37.8
Table 13: Average evaluation results of LLMs across different language versions on APEval.
Model C H, C C, U H, C, U Avg.
6B+ Models
DS-Coder-6.7B-Base 34.2 30.9 45.9 36.2 36.8
DS-Coder-6.7B-Inst 40.3 37.0 58.1 40.6 44.0
Yi-Coder-9B 32.1 37.0 28.9 35.8 33.4
Yi-Coder-9B-Chat 40.6 34.2 54.9 36.6 41.5
Qwen2.5-Coder-7B 49.6 40.7 62.2 39.8 48.1
Qwen2.5-Coder-7B-Inst 16.7 33.8 63.8 43.5 39.5
CursorCore-DS-6.7B 53.7 41.9 58.5 40.2 48.6
CursorCore-Yi-9B 51.7 45.5 64.6 41.9 50.9
CursorCore-QW2.5-7B 56.1 43.5 64.6 46.7 52.7
1B+ Models
DS-Coder-1.3B-Base 20.3 13.4 28.9 23.6 21.5
DS-Coder-1.3B-Inst 31.9 33.3 39.9 36.2 35.3
Yi-Coder-1.5B 16.7 19.1 13.8 22.0 17.9
Yi-Coder-1.5B-Chat 22.4 17.9 19.5 23.2 20.7
Qwen2.5-Coder-1.5B 36.2 31.3 46.8 39.9 38.5
Qwen2.5-Coder-1.5B-Inst 22.4 26.4 41.5 37.4 31.9
CursorCore-DS-1.3B 36.6 37.0 48.4 38.2 39.9
CursorCore-Yi-1.5B 41.8 35.3 50.8 37.4 41.1
CursorCore-QW2.5-1.5B 43.9 39.8 54.9 42.3 45.2
@@
Appendix M Chat template
Our model’s chat template (OpenAI, 2023) is adapted from the ChatML template, where each message in the conversation is restricted to one of the following roles: system, history, current, user, or assistant. The assistant’s output includes both code modifications and chat interaction with the user. To indicate code changes, we use two special tokens “<|next_start|>” and “<|next_end|>” to wrap the code modification parts. This approach models Assistant-Conversation effectively and is compatible with standard ChatML templates and chatbot applications. Figure 17 illustrates an example of our chat template, while Figure 18 presents examples of the chat template when using the LC and SR modes described in Appendix B.
@@
Appendix N Prompts for evaluation
@@
Appendix O Prompts for data collection
We design specific system prompts and few-shot examples to collect high-quality training data, as we find that many examples are very difficult to complete with current LLMs, and only a few of them can be successfully completed using rough prompts. For AI Programmer, we utilize LLMs to simulate programmers at three different skill levels, with each level using a distinct set of prompts as shown in Tables 14, 15 and 16. Additionally, prompts used for evaluating whether the outputs align with user intent, generating user instructions, and facilitating chat interactions between models and users are outlined in Tables 19, 17 and 18. Partial few-shot examples are shown in Figures 19, 20, 21, 22, 23 and 24.
Appendix P Limitations and future work
@@
In this work, we focus on supporting the development of single files or function-level code. However, real-world development operates at the repository level, involving multiple files and greater interaction with IDEs. Previous research has made notable advances in repository-level tasks such as code completion (Zhang et al., 2023), issue fixing (Jimenez et al., 2024), and documentation generation (Luo et al., 2024a). Repository-level code assistance deals with larger datasets, and achieving optimal performance and speed will require more effort. We leave the exploration of multi-file repository-level programming assistance and leveraging additional IDE interactions for future work.
@@
Our benchmark is relatively small and based on a multilingual extension of HumanEval, making it insufficient to cover all development scenarios. Beyond using the classic Pass@k metric to evaluate accuracy, other criteria should also be considered, such as evaluating the model’s efficiency, security, and redundancy (Huang et al., 2024; Pearce et al., 2021; Li et al., 2024a).
@@
Methods like PPO (Schulman et al., 2017) and DPO (Rafailov et al., 2023), which optimize models based on human preferences, have been widely used in LLMs. In programming assistance, programmers can provide feedback on predicted outputs for identical or similar coding processes, further optimizing the model (Shinn et al., 2023). To enable this, a significant amount of feedback data from programmers using AI-assisted tools should be collected or synthesized.
@@
We aim to integrate function calls (Patil et al., 2023) into the model to further enhance its capabilities. One potential application is incorporating function calls into the thinking process, such as retrieving information or executing partial code for feedback. Although our final models excludes this thinking step due to performance and speed considerations, we are exploring hybrid approaches to introduce this process while maintaining speed and combine it with other strategies for searching how to edit. Another application is leveraging function calls in output, where calling a Python script for tasks like variable replacement might be more efficient than manually generating code blocks or search-and-replace strategies. For repository-level changes, using terminal commands or IDE APIs could sometimes be a more convenient solution.